# Evaluating Large Language Models on Wikipedia-Style Survey Generation

### v3

@@ Abstract
Educational materials such as survey articles in specialized fields like computer science traditionally require tremendous expert inputs and are therefore expensive to create and update. Recently, Large Language Models (LLMs) have achieved significant success across various general tasks. However, their effectiveness and limitations in the education domain are yet to be fully explored. In this work, we examine the proficiency of LLMs in generating succinct survey articles specific to the niche field of NLP in computer science, focusing on a curated list of 99 topics. Automated benchmarks reveal that GPT-4 surpasses its predecessors like GPT-3.5, PaLM2, and LLaMa2 in comparison to the established ground truth. We compare both human and GPT-based evaluation scores and provide in-depth analysis. While our findings suggest that GPT-created surveys are more contemporary and accessible than human-authored ones, certain limitations were observed. Notably, GPT-4, despite often delivering outstanding content, occasionally exhibited lapses like missing details or factual errors. At last, we compared the rating behavior between humans and GPT-4 and found systematic bias in using GPT evaluation.
@@ 1 Introduction
Recently, large language models (LLMs) have attracted significant attention due to their strong performance on general natural language processing (NLP) tasks Shaib et al. (2023). Especially, the GPT family Brown et al. (2020) shows great ability in various applications. While it has been demonstrated that they perform well in many general tasks, their effectiveness in domain-specific tasks continues to be under scrutiny Tian et al. (2023). Specifically, the text produced by LLMs can sometimes exhibit issues like creating false information and hallucination Zhao et al. (2023).
In the context of scientific education, automatic survey generation aims to employ machine learning or NLP techniques to create a structured overview of a specific concept Sun and Zhuge (2022); Li et al. (2022). Automating this process not only alleviates the manual effort but also ensures timely updates at a reduced cost. A common approach involves an initial information retrieval phase to select pertinent documents or sentences based on the query topic. This is followed by a summarization or simplification phase to produce the final survey Jha et al. (2013); Li et al. (2022). While LLMs have the potential to be an alternative method for writing scientific surveys, their effectiveness and limitations are not yet thoroughly investigated.
@@ 2 Method
We adopt the Surfer100 dataset Li et al. (2022), which contains 100 manually written survey articles on NLP concepts. Each survey contains five sections: Introduction, History, Key Ideas, Uses/Applications, and Variations. Each section includes 50-150 tokens. Then we compare three settings: zero-shot (ZS), one-shot (OS) and description prompt (DP). For zero-shot, we directly ask the model to generate the article by providing the following prompt: Generate a survey about <Topic>. There should be five sub-sections: Introduction, History, Key Ideas, Variations and Applications. Each subsection should contain 50-150 words. For the one-shot setting, we add a sample survey article; for the description prompt setting, we add a detailed description to each section explaining what should be included. For example, SECTION 1: INTRODUCTION Describe what the topic is (a method, a model, a task, a dataset), which field/subfield it is part of, quick overview of applications and motivation behind the concept and related ideas). To further enrich the provided information, we also introduce a combination of one-shot and description prompt (OSP). The full prompt is shown in Appendix A. By employing a single ground truth for one-shot learning, we generate 99 surveys per setting. Moreover, we evaluate a special retrieval augmented generation (RAG) Gao et al. (2023); Ram et al. (2023); Yang et al. (2023); Aksitov et al. (2023) setting (OS+IR) which links to Wikipedia articles and web data.
@@
Automatic Evaluation We evaluate the generated surveys using a range of automatic metrics including ROUGE, BERTScore Zhang* et al. (2020), MoverScore Zhao et al. (2019), UniEval Zhong et al. (2022) and BARTScore Yuan et al. (2021). Tab. 1 provides an overview of results for the following LLMs: LLaMa2 (13B, 70B), PaLM2 (text-bison), GPT-3.5 (Turbo-0613) as well as GPT-4 (0613) across different prompt settings. We first notice that GPT-4 consistently outperforms other baselines, obtaining a significant improvement of around 2% to 20% when enhancing prompts. Specifically, GPT-4 OSP achieves the top spot under most situations. However, it is not to say that prompt enrichment always yields positive results. For instance, in the case of LLaMa2, one-shot and description prompts perform better than OSP. As for PaLM2, four types of prompts obtain similar results. When we add external knowledge (GPT-4 OS+IR), there is some improvement compared to GPT-4 OS. As our primary goal is to study the extent of knowledge LLMs possess in this task, we mainly focus on analyses in settings without external data. However, additional analysis about GPT-4 OS+IR setting can be found in Appendix C.
Human and GPTs Evaluation We employ two NLP experts, GPT-4 and G-Eval Liu et al. (2023a) to evaluate surveys generated by the best GPT-4 OSP setting, focusing on 6 perspectives: Readability, Relevancy, Hallucination, Completeness, Factuality. The detailed guidelines can be found in Appendix B. It’s important to note that we implement a pre-selection stage in the choice of human experts (Appendix B). Tab. 2 shows that both human experts and GPTs agree that the generated surveys perform well across most aspects, though the completeness exhibits marginally lowest scores. According to IAA, we can observe that human experts demonstrate a high consistent quality of the generated surveys whlie GPT-4 and G-Eval have more randomness. To better understand the degree of agreement between human experts and GPT-4 on ratings, we also calculate Kendall’s and -value as shown in Tab. 3. We can observe that the Factuality possesses the highest degree of correlation. In contrast, Redundancy displays the lowest correlation while the other aspects exhibit relatively lower correlation levels. This difference is largely because Factuality is based on objective ground truth, while Redundancy is more dependent on subjective judgment. Notably, we can conclude that in most scenarios, GPT-4 showcases similar evaluative opinions as humans, despite showing a higher degree of variability across different independent sessions. Regarding RQ1 and RQ2, we find that 1) LLMs can produce high-quality survey articles, and 2) with specific guidance, there’s a strong consistency between GPT outputs and human judgment.
@@ 3 Analysis
Novel Entity Mention To further investigate how interesting the generated content is, we look at the mentions of novel entities following Lee et al. (2022). Specifically, we examine the survey content, juxtaposing the entities within it against those in the ground truth. We employed Stanza Qi et al. (2020) to identify all entities among all types present in both the LLM-generated text and the ground truth. Subsequently, we quantified the number of unique entities that emerged in the LLM-generated content. For an equitable comparison, we assessed the one-shot with prompt settings of LLaMa2-13b, PaLM2, and GPT-4, in addition to the ZS setting of GPT-3.5, as depicted in Fig. 2. Our findings reveal that PaLM2 exhibited the least variation in entity mentions, while LLaMa2-13b showcased the most. Despite GPT-4’s outstanding performance in both automated and human evaluations, we didn’t discern a marked novelty in its entity mentions. We speculate that this might be an inherent compromise when generating high-fidelity content in relation to the ground truth. So far, regarding RQ1, although LLMs register commendable results based on predefined criteria, certain shortcomings are evident. Specifically, we observe some omitted details, particularly within the Introduction and History sections. While LLMs often introduce new entities, we don’t find a significant correlation between this tendency and their performance. More case studies are in Appendix D.
@@ 4 Discussion and Conclusion
Limitations
While GPT-4 can produce contemporary and accessible content, there are instances where the depth and detailing of topics are compromised, leading to potential information gaps. GPT-4 occasionally produces factual inaccuracies in the content. This necessitates additional verification, especially for educational materials where accuracy is paramount. When comparing ratings between human experts and GPT-4, a systematic bias was observed in GPT evaluations. This can skew outcomes and potentially mislead the quality perception of the generated content.
Ethical Considerations
@@
Appendix A Prompt Example
Appendix B Human Evaluation Guidance
@@
Appendix C Comparisons with External Knowledge
@@
Appendix D More Case Study and Observations
D.1 Understanding of “Survey”
@@ D.2 Incomplete Information
D.3 Nebulous Sentence Structure
@@
D.4 High-quality Survey
@@
D.5 Going Beyond the Ground Truth
@@
Appendix E Potential Risks
@@
Appendix F Experimental Setup
In our work, we mainly utilize the paid GPT-4 model to generate Wiki-style survey articles and further explore its capabilities to score the top-performing predictions compared to references. The total cost of these experiments is around 230 USD. During the automatic evaluation stage, we compute the ROUGE score and BERTScore using the officially provided APIs: rouge333https://pypi.org/project/rouge/ and bert_score444https://github.com/Tiiiger/bert_score. For calculations involving MoverScore, UniEval, and BARTScore, we directly download their source codes. All experiments were performed using the high-performance machine with 4 A100 40GB NVIDIA cards. As the experiments do not involve fine-tuning, for each setting, we were able to finish in a few hours. As for the human evaluation stage, we calculate the Krippendorff’s and Kendall’s scores with the authorized APIs krippendorff555https://pypi.org/project/krippendorff/ and scipy666https://scipy.org/.

### v4

@@ Abstract
Educational materials such as survey articles in specialized fields like computer science traditionally require tremendous expert inputs and are therefore expensive to create and update. Recently, Large Language Models (LLMs) have achieved significant success across various general tasks. However, their effectiveness and limitations in the education domain are yet to be fully explored. In this work, we examine the proficiency of LLMs in generating succinct survey articles specific to the niche field of NLP in computer science, focusing on a curated list of 99 topics. Automated benchmarks reveal that GPT-4 surpasses its predecessors, inluding GPT-3.5, PaLM2, and LLaMa2 by margins ranging from 2% to 20% in comparison to the established ground truth. We compare both human and GPT-based evaluation scores and provide in-depth analysis. While our findings suggest that GPT-created surveys are more contemporary and accessible than human-authored ones, certain limitations were observed. Notably, GPT-4, despite often delivering outstanding content, occasionally exhibited lapses like missing details or factual errors. At last, we compared the rating behavior between humans and GPT-4 and found systematic bias in using GPT evaluation.
@@ 1 Introduction
Recently, large language models (LLMs) have attracted significant attention due to their strong performance on general natural language processing (NLP) tasks Shaib et al. (2023); Feng et al. (2022). Especially, the GPT family Brown et al. (2020) shows great ability in various applications. While it has been demonstrated that they perform well in many general tasks, their effectiveness in domain-specific tasks continues to be under scrutiny Tian et al. (2023). Specifically, the text produced by LLMs can sometimes exhibit issues like creating false information and hallucination Zhao et al. (2023); Yang et al. (2024a).
In the context of scientific education, automatic survey generation aims to employ machine learning or NLP techniques to create a structured overview of a specific concept Sun and Zhuge (2022); Li et al. (2022); Yang et al. (2024b). Automating this process not only alleviates the manual effort but also ensures timely updates at a reduced cost. A common approach involves an initial information retrieval phase to select pertinent documents or sentences based on the query topic. This is followed by a summarization or simplification phase to produce the final survey Jha et al. (2013); Li et al. (2022). While LLMs have the potential to be an alternative method for writing scientific surveys, their effectiveness and limitations are not yet thoroughly investigated.
@@ 2 Method
2.1 Dataset
We adopt the Surfer100 dataset Li et al. (2022), which contains 100 manually written Wikipedia-style articles on NLP concepts. Each survey is structured into five sections: Introduction, History, Key Ideas, Uses/Applications, and Variations, with each section containing between 50 and 150 words.
2.2 Experiment Setup
We compare three settings: zero-shot (ZS), one-shot (OS) and description prompt (DP). For zero-shot, we directly ask the model to generate the article by providing the following prompt: Generate a survey about <Topic>. There should be five sub-sections: Introduction, History, Key Ideas, Variations and Applications. Each subsection should contain 50-150 words. For the one-shot setting, we add a ground-truth article of Word2Vec as the sample survey; for the description prompt setting, we add a detailed description to each section explaining what should be included. For example, SECTION 1: INTRODUCTION Describe what the topic is (a method, a model, a task, a dataset), which field/subfield it is part of, quick overview of applications and motivation behind the concept and related ideas). To further enrich the provided information, we also introduce a combination of one-shot and description prompt (OSP). The full prompt is shown in Figure 1. By employing a single ground truth for one-shot learning, we generate 99 surveys per setting. Moreover, we evaluate a special retrieval augmented generation (RAG) Gao et al. (2023); Ram et al. (2023); Yang et al. (2023); Aksitov et al. (2023) setting (denoted as OS+IR) which enables GPT-4 link to Wikipedia pages and access to web data. We elaborate more experimental details in Appendix B
@@ 2.2 Experiment Setup
2.3 Evaluation Metrics
Automatic Evaluation We evaluate the generated surveys using a range of automatic metrics including ROUGE, BERTScore Zhang* et al. (2020), MoverScore Zhao et al. (2019), UniEval Zhong et al. (2022) and BARTScore Yuan et al. (2021). Tab. 1 provides an overview of results for the following LLMs: LLaMa2 (13B, 70B), PaLM2 (text-bison), GPT-3.5 (Turbo-0613) as well as GPT-4 (0613) across different prompt settings. We first notice that GPT-4 consistently outperforms other baselines, obtaining a significant improvement of around 2% to 20% when enhancing prompts. Specifically, GPT-4 OSP achieves the top spot under most situations. However, it is not to say that prompt enrichment always yields positive results. For instance, in the case of LLaMa2, one-shot and description prompts perform better than OSP. As for PaLM2, four types of prompts obtain similar results. When we add external knowledge (GPT-4 OS+IR), there is some improvement compared to GPT-4 OS. As our primary goal is to study the extent of knowledge LLMs possess in this task, we mainly focus on analyses in settings without external data. However, additional analysis about GPT-4 OS+IR setting can be found in Appendix B.
Human and GPTs Evaluation We employ two NLP experts, GPT-4 and G-Eval Liu et al. (2023a) to evaluate surveys generated by the best GPT-4 OSP setting, focusing on 6 perspectives: Readability, Relevancy, Hallucination, Completeness, Factuality. Both GPT models and humans are required to score each aspect on a scale from to , following the same guidelines. The detailed guidance can be found in Appendix A. It’s important to note that we implement a pre-selection stage in the choice of human experts (Appendix A). Tab. 2 shows that both human experts and GPTs agree that the generated surveys perform well across most aspects, though the completeness exhibits marginally lowest scores. According to IAA, we can observe that human experts demonstrate a high consistent quality of the generated surveys whlie GPT-4 and G-Eval have more randomness. To better understand the degree of agreement between human experts and GPT-4 on ratings, we also calculate Kendall’s and -value as shown in Tab. 3. We can observe that the Factuality possesses the highest degree of correlation. In contrast, Redundancy displays the lowest correlation while the other aspects exhibit relatively lower correlation levels. This difference is largely because Factuality is based on objective ground truth, while Redundancy is more dependent on subjective judgment. Notably, we can conclude that in most scenarios, GPT-4 showcases similar evaluative opinions as humans, despite showing a higher degree of variability across different independent sessions. Regarding RQ1 and RQ2, we find that 1) LLMs can produce high-quality survey articles, and 2) with specific guidance, there’s a strong consistency between GPT outputs and human judgment.
@@ 3 Analysis
Novel Entity Mention To further investigate how interesting the generated content is, we look at the mentions of novel entities following Lee et al. (2022). Specifically, we examine the survey content by comparing the entities it contains with those in the ground truth. We employ Stanza Qi et al. (2020) to identify all entities in both the LLM-generated text and the ground truth. Subsequently, we quantify the number of unique entities found in the LLM-generated content. For a fair comparison, we analyze the one-shot with prompt settings of LLaMa2-13b, PaLM2, and GPT-4, in addition to the ZS setting of GPT-3.5, as depicted in Fig. 3. Our findings reveal that PaLM2 exhibited the least variation in entity mentions, while LLaMa2-13b showcased the most. Despite GPT-4’s outstanding performance in both automated and human evaluations, we didn’t discern a marked novelty in its entity mentions. We speculate that this might be an inherent compromise when generating high-fidelity content in relation to the ground truth. So far, regarding RQ1, although LLMs register commendable results based on predefined criteria, certain shortcomings are evident. Specifically, we observe some omitted details, particularly within the Introduction and History sections. While LLMs often introduce new entities, we don’t find a significant correlation between this tendency and their performance. More case studies are in Appendix C.
@@ 4 Discussion and Conclusion
Acknowledgements
This work was partially supported by the Deutsche Forschungsgemeinschaft (DFG) through the priority program RATIO (SPP-1999), Grant No. 376059226. It was also supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI, Grant Number 24K20832. We would like to thank these funding agencies for their generous support. Additionally, we would like to thank the anonymous reviewers for their helpful feedback.
Limitations and Ethical Considerations
GPT-4 can generate contemporary, accessible content, but sometimes compromises depth and detail, leading to potential information gaps and occasional factual inaccuracies. This requires extra verification. Comparing ratings between human experts and GPT-4 revealed a systematic bias in GPT evaluations, which can skew outcomes and mislead quality perception.
@@ Acknowledgements
Appendix A Human Evaluation Guidance
@@
Appendix B Comparisons with External Knowledge
@@
Appendix C More Case Study and Observations
C.1 Understanding of “Survey”
@@ C.2 Incomplete Information
C.3 Nebulous Sentence Structure
@@
C.4 High-quality Survey
@@
C.5 Going Beyond the Ground Truth
@@
Appendix D Potential Risks
@@
Appendix E Experimental Details
In our work, we mainly utilize the paid GPT-4 model to generate Wiki-style survey articles and further explore its capabilities to score the generations. The total cost of these experiments is around 230 USD. During the automatic evaluation stage, we compute the ROUGE score and BERTScore using the officially provided APIs: rouge333https://pypi.org/project/rouge/ and bert_score444https://github.com/Tiiiger/bert_score. For calculations involving MoverScore, UniEval, and BARTScore, we directly download their source codes. All experiments were performed using the high-performance machine with 4 A100 40GB NVIDIA cards. As the experiments do not involve fine-tuning, for each setting, we were able to finish in a few hours. As for the human evaluation stage, we calculate the Krippendorff’s and Kendall’s scores with the authorized APIs krippendorff555https://pypi.org/project/krippendorff/ and scipy666https://scipy.org/.